bruce is 38 years old, his son is 8 years old. In how many years will bruce be 3 times older than his son?
Answer by Maths68(1474)

Let after x years Bruce will be three times older then his son
After x years Bruce will be = 38+x
After x years Son will be = 8+x
Then
Bruce will be three times as old as his son
38+x=3(8+x)
38+x=24+3x
x-3x=24-38
-2x=-14
-2x/-2=-14/-2
x=7

After 7 years Bruce will be three times older then his son

After 7 years Bruce will be = 38+x = 38+7 = 45
After 7 years Son will be = 8+x = 8+7 = 15